OpenStudy (anonymous):

calculate the problem 1/3 and 1/4

OpenStudy (anonymous):

you want the answer of this question: 1/3 = ? and 1/4 = ? ?!

OpenStudy (anonymous):

at the end of the ? is multiple choice a=7/12, b=1, c=4/3, d=1/12